The following data were collected related to the number of visits to a suburban health clinic.

Month Vists
January 1610
February 1585
March 1649
April 1590
May 1540
June 1397
July 1410
August 1350
September 1495
Ocotober 1564
November 1602
December 1655

The null hypothesis is that the number of visits is uniformly distributed throughout the year. Use a chi-square goodness-of-fit test to determine whether the null hypothesis should be rejected at α = 0.05. Interpret the results.
Answer #1
null hypothesis:Ho:number of visits are uniformly distributed.through out the year
Alternate hypothesis:Ho:number of visits are not uniformly distributed.through out the year
degree of freedom =categories-1= 11
for 0.05 level and 11 df :crtiical value X2 = 19.675
Decision rule: reject Ho if value of test statistic X2>19.675
applying chi square goodness of fit test:
           relative observed Expected residual Chi square
category frequency(p) Oi Ei=total*p R2i=(Oi-Ei)/√Ei R2i=(Oi-Ei)2/Ei
Jan    1/12 1610.000 1537.25 1.86 3.443
Feb    1/12 1585.000 1537.25 1.22 1.483
Mar    1/12 1649.000 1537.25 2.85 8.124
Apr    1/12 1590.000 1537.25 1.35 1.810
May    1/12 1540.000 1537.25 0.07 0.005
Jun    1/12 1397.000 1537.25 -3.58 12.796
Jul    1/12 1410.000 1537.25 -3.25 10.533
Aug    1/12 1350.000 1537.25 -4.78 22.809
Sep    1/12 1495.000 1537.25 -1.08 1.161
Oct    1/12 1564.000 1537.25 0.68 0.465
Nov    1/12 1602.000 1537.25 1.65 2.727
Dec    1/12 1655.000 1537.25 3.00 9.019
total 1.000 18447 18447 74.3758
test statistic X2 = 74.376
since test statistic falls in rejection region we reject null hypothesis
we have sufficient evidence to conclude that number of visits are not uniformly distributed.through out the year